OpenStudy (pinklion23):

Find the slope y+7=-3/4(x+4)

OpenStudy (campbell_st):

well you have an equation in point slope form \[y - y_{1} = m(x - x_{1})\] m is the slope... match it to your equation, to get the answer

OpenStudy (pinklion23):

ok thank you oh but after i find the answer can you show me all of the steps to find the answer and the answer is -3/4

OpenStudy (campbell_st):

there are no steps... its a case of just matching in your question \[y_{1} = -7~~~x_{1} = -4\] so the point was (-7, -4) and a slope of -3/4

OpenStudy (campbell_st):

the equation is \[y - (-7) = -3/4(x - (-4))\] oops... the point should be (-4, -7) x 1st

OpenStudy (pinklion23):

ok thank you